# Line Defects with a Cusp in Fermionic CFTs

### v1

@@ 1 Introduction and Summary
Figure 1.1: Cusp at angle $\theta$ with two different defect couplings on each arm
@@
Figure 1.2: The two defects $a$ and $b$ with cusp angle $\theta\rightarrow 0$ can be described as two line defects $a$ and $\bar{b}$ fusing
@@ 2.1 Cusp Anomalous dimension in the GNY model
Figure 2.2: Diagrams used to compute the self interaction energy to $\mathcal{O}(\epsilon)$.
@@
Note that the coefficient of $\log(M^{2})$ in this expression is proportional to $\beta_{h_{1}}/h_{1}+\beta_{h_{2}}/h_{2}$, where $\beta_{h}$ is the beta function of the defect coupling given in eq. (B). Hence, at the fixed point the dependence on $M$ drops out as expected.
@@
$\begin{split}\Gamma_{h_{1}h_{2}}(\theta)=&-\frac{1}{4\pi^{2}}\left(h_{1*}h_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(h_{1*}^{2}+h_{2*}^{2})}{2}\right)\\ &-\frac{1}{4\pi^{2}}\left(h_{1*}h_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(h_{1*}^{2}+h_{2*}^{2})}{2}\right)\Big(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})\Big)\\ &+\frac{g_{2*}}{384\pi^{4}}\log(64\pi e^{\gamma-1})\left(h_{1*}h_{2*}\frac{(h_{1*}^{2}+h_{2*}^{2})}{2}\frac{\pi-\theta}{\sin\theta}-\frac{(h_{1*}^{4}+h_{2*}^{4})}{2}\right)\\ &-\left(\epsilon\frac{1}{8\pi^{2}}-\frac{g_{1*}^{2}N}{128\pi^{4}}\right)(h_{1*}h_{2*}F^{\prime}_{0}(\cos\theta)-(h_{1*}^{2}+h_{2*}^{2}))\\ &+\frac{g_{2*}}{1536\pi^{8}}((h_{1*}^{3}h_{2*}+h_{1*}h_{2*}^{3})I_{11}(\cos\theta)-(8\pi^{4}-4\pi^{4}\log 8-24\pi^{2}\zeta(3))(h_{1*}^{4}+h_{2*}^{4}))\\ &+\frac{g_{2*}}{1024\pi^{8}}((h_{1*}^{2}h_{2*}^{2})I_{12}(\cos\theta)-12\pi^{2}\zeta(3)(h_{1*}^{4}+h_{2*}^{4}))\end{split}$ (2.28)
@@
$\begin{split}\Gamma_{h_{1}h_{2}}(\theta)=&-\frac{h_{*}^{2}}{4\pi^{2}}\left(\text{sgn}(h_{1}h_{2})\frac{\pi-\theta}{\sin\theta}-1\right)\\ &-\frac{h_{*}^{2}}{4\pi^{2}}\left(\text{sgn}(h_{1}h_{2})\frac{\pi-\theta}{\sin\theta}-1\right)\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)(\text{sgn}(h_{1}h_{2})F^{\prime}_{0}(\cos\theta)-2)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}(\text{sgn}(h_{1}h_{2})I_{11}(\cos\theta)-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi^{8}}(I_{12}(\cos\theta)-24\pi^{2}\zeta(3))\end{split}$ (2.29)
@@ 2.2 Generalizing to $N_{s}$ scalar fields
$S_{N,N_{s}}=\int d^{d}x\left[\frac{1}{2}(\partial\mathbf{\phi})^{2}+\bar{\Psi}\not{\partial}\Psi+\frac{g_{2}}{24}(\mathbf{\phi}\cdot\mathbf{\phi})^{2}+g_{1}\sum_{a}\phi_{a}\bar{\Psi}\Sigma_{a}\Psi\right]$ (2.34)
@@
$\Delta_{\phi}=1+\frac{\partial\beta_{h}}{\partial h}\Bigg|_{h=0}=1-\frac{\epsilon}{2}+\frac{g_{1}^{2}N}{2(4\pi)^{2}}+\frac{g_{2}^{2}(2+N_{s})}{36(4\pi)^{4}}-\frac{g_{1}^{4}N(4+N_{s})}{4(4\pi)^{4}}$ (2.43)
@@
$\Delta_{\hat{\phi}}$= $1-\frac{\epsilon}{2}+\frac{1}{(4\pi)^{2}}\left(\frac{g_{1}^{2}N}{2}+\frac{g_{2}h_{*}^{2}}{2}\right)+\frac{1}{(4\pi)^{4}}\Bigg\{g_{2}^{2}\left(\frac{(2+N_{s})}{36}-\frac{h_{*}^{2}(N_{s}+8)}{12}-\frac{5h_{*}^{4}}{12}\right)-\frac{3g_{2}g_{1}^{2}h_{*}^{2}N}{4}$
$+g^{4}_{1}\left(-\frac{(N_{s}+4)N}{4}+3h_{*}^{2}N\left(1-\frac{\pi^{2}}{6}\right)\right)\Bigg\}$ (2.44)
@@
$\begin{split}\Gamma_{\mathbf{h}_{1*}\mathbf{h}_{2*}}(\theta)=&-\frac{1}{4\pi^{2}}\left(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})}{2}\right)\\ &-\frac{1}{4\pi^{2}}\left(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})}{2}\right)\Big(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})\Big)\\ &+\frac{g_{2*}}{384\pi^{4}}\log(64\pi e^{\gamma-1})\left(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*}\frac{(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})}{2}\frac{\pi-\theta}{\sin\theta}-\frac{((\mathbf{h}_{1*}^{2})^{2}+(\mathbf{h}_{2*}^{2})^{2})}{2}\right)\\ &-\left(\epsilon\frac{1}{8\pi^{2}}-\frac{g_{1*}^{2}N}{128\pi^{4}}\right)((\mathbf{h}_{1*}\cdot\mathbf{h}_{2*})F^{\prime}_{0}(\cos\theta)-(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2}))\\ &+\frac{g_{2*}}{1536\pi^{8}}\Big((\mathbf{h}_{1*}\cdot\mathbf{h}_{2*})(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})I_{11}(\cos\theta)-(8\pi^{4}-4\pi^{4}\log 8-24\pi^{2}\zeta(3))((\mathbf{h}_{1*}^{2})^{2}+(\mathbf{h}_{2*}^{2})^{2})\Big)\\ &+\frac{g_{2*}}{1024\pi^{8}}\Big(\left(\frac{\mathbf{h}_{1*}^{2}\mathbf{h}_{2*}^{2}+2(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*})^{2}}{3}\right)I_{12}(\cos\theta)-12\pi^{2}\zeta(3)\big((\mathbf{h}_{1*}^{2})^{2}+(\mathbf{h}_{2*}^{2})^{2}\big)\Big)\end{split}$ (2.46)
@@
$\begin{split}\Gamma_{\mathbf{h}_{1}\mathbf{h}_{2}}(\theta)=&-\frac{h_{*}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\\ &-\frac{h_{*}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)((\hat{m}_{1}\cdot\hat{m}_{2})F^{\prime}_{0}(\cos\theta)-2)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}((\hat{m}_{1}\cdot\hat{m}_{2})I_{11}(\cos\theta)-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi^{8}}(\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)I_{12}(\cos\theta)-24\pi^{2}\zeta(3))\end{split}$ (2.47)
@@
$\begin{split}\Gamma_{1}(\theta)=&-\frac{h_{*,1}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\\ &-\frac{h_{*,0}^{2}}{4\pi^{2}}\Bigg[-\frac{\epsilon(N-4+N_{s}+(4-N_{s})\log 64)}{N+8-2N_{s}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\\ &+\Big(\frac{\epsilon(8-2N_{s})}{2(N+8-2N_{s})}(\hat{m}_{1}\cdot\hat{m}_{2}F^{\prime}_{0}(\cos\theta)-2)\Big)\\ &-\frac{3\epsilon(4-N_{s})}{2\pi^{4}(N+8-2N_{s})}\times\\ &\Big(\frac{1}{3}((\hat{m}_{1}\cdot\hat{m}_{2})I_{11}(\cos\theta)-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &+\frac{1}{4}(\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)I_{12}(\cos\theta)-24\pi^{2}\zeta(3))\Big)\Bigg]\end{split}$ (2.50)
@@ 2.3 Large $N$ limit
Figure 2.3: Large NN cusp diagram.
@@ 3 Results and Discussion
Figure 3.1: The term of $\mathcal{O}(\epsilon)$ of the cusp anomalous dimension in the two types of cusped defect in the GNY model
@@
Figure 3.2: The term of $\mathcal{O}(\epsilon^{0})$ of the cusp anomalous dimension in the two types of cusped defect in the GNY model
Figure 3.3: Padé Approximation of $\Gamma_{++}(\theta)$ in the $d=3$ GNY model
@@
$\Delta_{\hat{\phi}}$ (Padé(1,1) using a 2-loop calculation) 1.3673 1.3386 1.3076
@@ 3.1 Dimension of Defect Changing and Creation (Annihilation) Operators
$\begin{split}\Delta_{\hat{m}_{1},\hat{m}_{2}}=&-\frac{h_{*}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}-1\right)\\ &-\frac{h_{*}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}-1\right)\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{4\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{64\pi^{2}}\right)((\hat{m}_{1}\cdot\hat{m}_{2})-1)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}((\hat{m}_{1}\cdot\hat{m}_{2})-1)(8\pi^{4}-4\pi^{4}\log 8-24\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi^{8}}(\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)-1)(24\pi^{2}\zeta(3))\end{split}$ (3.1)
@@
$\begin{split}\Delta_{\hat{m}_{1},\hat{m}_{2}}=&(1-\hat{m}_{1}\cdot\hat{m}_{2})\Big[\frac{h_{*}^{2}}{4\pi^{2}}\\ &+\frac{h_{*}^{2}}{4\pi^{2}}\left(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\right)\\ &+\left(\epsilon\frac{h_{*}^{2}}{4\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{64\pi^{4}}\right)\\ &-\frac{g_{2*}h_{*}^{4}}{192\pi^{8}}(2\pi^{4}-\pi^{4}\log 8-6\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}\zeta(3)}{64\pi^{6}}(1+\hat{m}_{1}\cdot\hat{m}_{2})\Big]\end{split}$ (3.2)
@@
$\begin{split}\Delta_{\hat{m}_{1}\hat{m}_{2}}^{(1)}=&(1-\hat{m}_{1}\cdot\hat{m}_{2})\Bigg[\frac{h_{*,1}^{2}}{4\pi^{2}}\\ &+\frac{h_{*,0}^{2}}{4\pi^{2}}\Bigg(-\frac{(N-4+N_{s}+(4-N_{s})\log 64)}{N+8-2N_{s}}\\ &+\frac{(8-2N_{s})}{(N+8-2N_{s})}\\ &-\frac{3(4-N_{s})}{2\pi^{4}(N+8-2N_{s})}\times\\ &\Big(\frac{8\pi^{4}-4\pi^{4}\log 8-24\pi^{2}\zeta(3)}{3}+16\pi^{2}\zeta(3)(1+\hat{m}_{1}\cdot\hat{m}_{2})\Big)\Bigg)\Bigg]\end{split}$ (3.5)
@@
We obtain
$\Delta_{0,\mathbf{h}}^{(0)}=\frac{h_{*,0}^{2}}{8\pi^{2}}$ (3.9)
and
$\begin{split}\Delta_{0,\mathbf{h}}^{(1)}=&\frac{h_{*,1}^{2}}{8\pi^{2}}\\ &+\frac{h_{*,0}^{2}}{8\pi^{2}}\Big(\frac{\epsilon}{2}\log(\pi e^{\gamma})-\frac{g_{1*}^{2}N}{32\pi^{2}}\log(\pi e^{\gamma+2})-\frac{g_{2*}h_{*,0}^{2}}{96\pi^{2}}\log(64\pi e^{\gamma-1})\Big)\\ &+\left(\epsilon\frac{h_{*,0}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*,0}^{2}}{128\pi^{4}}\right)\\ &-\frac{g_{2*}h_{*,0}^{4}}{384\pi^{6}}\Big(2\pi^{2}-\pi^{2}\log 8-6\zeta(3)\Big)\\ &-\frac{3g_{2*}h_{*,0}^{4}}{256\pi^{6}}\zeta(3)\end{split}$ (3.10)
@@
$\Delta_{0,\mathbf{h}}^{(0)}=\frac{(4-N_{s})(N_{s}+8)}{2\pi^{2}\Big(8-N-2N_{s}+\sqrt{N^{2}+4(-4+N_{s})^{2}+4N(28+5N_{s})}\Big)}$ (3.11)
$\begin{split}\Delta_{0,\mathbf{h}}^{(1)}=\frac{h_{*,1}^{2}}{8\pi^{2}}+\frac{h_{*,0}^{2}}{8\pi^{2}}\Bigg(&-\frac{\epsilon(N+8-2N_{s}+(4-N_{s})\log 64)}{N+8-2N_{s}}\\ &+\frac{\epsilon(8-2N_{s})}{(N+8-2N_{s})}\\ &-\frac{\epsilon(4-N_{s})}{\pi^{2}(N+8-2N_{s})}\Big(4\pi^{2}-2\pi^{2}\log 8-3\zeta(3)\Big)\Bigg)\end{split}$ (3.12)

### v2

@@ 1 Introduction and Summary
Figure 1.1: Cusp at angle $\theta$ with two different defect couplings on each arm.
@@
Figure 1.2: The two defects $a$ and $b$ with cusp angle $\theta\rightarrow 0$ can be described as two line defects $a$ and $\bar{b}$ fusing.
@@ 2.1 Cusp Anomalous dimension in the GNY model
Figure 2.2: Self-energy contributions to $\mathcal{O}(\epsilon)$.
@@
Note that the coefficient of $\log(\pi e^{\gamma}M^{2})$ in this expression is proportional to $\beta_{h_{1}}/h_{1}+\beta_{h_{2}}/h_{2}$, where $\beta_{h}$ is the beta function of the defect coupling given in eq. (B). Hence, at the fixed point the dependence on $M$ drops out as expected.
@@
$\begin{split}\Gamma_{h_{1}h_{2}}(\theta)=&-\frac{1}{4\pi^{2}}\left(h_{1*}h_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(h_{1*}^{2}+h_{2*}^{2})}{2}\right)\\ &+\frac{g_{1*}^{2}N}{64\pi^{4}}\left(h_{1*}h_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(h_{1*}^{2}+h_{2*}^{2})}{2}\right)\\ &+\frac{g_{2*}}{384\pi^{4}}\log(64e^{-1})\left(h_{1*}h_{2*}\frac{(h_{1*}^{2}+h_{2*}^{2})}{2}\frac{\pi-\theta}{\sin\theta}-\frac{(h_{1*}^{4}+h_{2*}^{4})}{2}\right)\\ &-\left(\epsilon\frac{1}{8\pi^{2}}-\frac{g_{1*}^{2}N}{128\pi^{4}}\right)(h_{1*}h_{2*}F^{\prime}_{0}(\cos\theta)-(h_{1*}^{2}+h_{2*}^{2}))\\ &+\frac{g_{2*}}{1536\pi^{8}}((h_{1*}^{3}h_{2*}+h_{1*}h_{2*}^{3})I_{11}(\cos\theta)-(8\pi^{4}-4\pi^{4}\log 8-24\pi^{2}\zeta(3))(h_{1*}^{4}+h_{2*}^{4}))\\ &+\frac{g_{2*}}{1024\pi^{8}}((h_{1*}^{2}h_{2*}^{2})I_{12}(\cos\theta)-12\pi^{2}\zeta(3)(h_{1*}^{4}+h_{2*}^{4}))\end{split}$ (2.28)
@@
$\begin{split}\Gamma_{h_{1}h_{2}}(\theta)=&-\frac{h_{*}^{2}}{4\pi^{2}}\left(\text{sgn}(h_{1}h_{2})\frac{\pi-\theta}{\sin\theta}-1\right)\\ &+\frac{h_{*}^{2}}{64\pi^{4}}\left(\text{sgn}(h_{1}h_{2})\frac{\pi-\theta}{\sin\theta}-1\right)\left(g_{1*}^{2}N+\frac{g_{2*}h_{*}^{2}}{6}\log(64e^{-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)(\text{sgn}(h_{1}h_{2})F^{\prime}_{0}(\cos\theta)-2)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}(\text{sgn}(h_{1}h_{2})I_{11}(\cos\theta)-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi^{8}}(I_{12}(\cos\theta)-24\pi^{2}\zeta(3))\end{split}$ (2.29)
@@ 2.2 Generalizing to $N_{s}$ scalar fields
$S_{N,N_{s}}=\int d^{d}x\left[\frac{1}{2}(\partial\mathbf{\phi})^{2}-\bar{\Psi}\not{\partial}\Psi+\frac{g_{2}}{24}(\mathbf{\phi}\cdot\mathbf{\phi})^{2}+g_{1}\sum_{a}\phi_{a}\bar{\Psi}\Sigma_{a}\Psi\right]$ (2.34)
@@
$\Delta_{\phi}=1+\frac{\partial\beta_{h}}{\partial h}\Bigg|_{h=0}=1-\frac{\epsilon}{2}+\frac{g_{1*}^{2}N}{2(4\pi)^{2}}+\frac{g_{2*}^{2}(2+N_{s})}{36(4\pi)^{4}}-\frac{g_{1*}^{4}N(4+N_{s})}{4(4\pi)^{4}}$ (2.43)
@@
$\Delta_{\hat{\phi}}$= $1-\frac{\epsilon}{2}+\frac{1}{(4\pi)^{2}}\left(\frac{g_{1*}^{2}N}{2}+\frac{g_{2*}h_{*}^{2}}{2}\right)+\frac{1}{(4\pi)^{4}}\Bigg\{g_{2*}^{2}\left(\frac{(2+N_{s})}{36}-\frac{h_{*}^{2}(N_{s}+8)}{12}-\frac{5h_{*}^{4}}{12}\right)-\frac{3g_{2*}g_{1*}^{2}h_{*}^{2}N}{4}$
$+g^{4}_{1*}\left(-\frac{(N_{s}+4)N}{4}+3h_{*}^{2}N\left(1-\frac{\pi^{2}}{6}\right)\right)\Bigg\}$ (2.44)
@@
$\begin{split}\Gamma_{\mathbf{h}_{1*}\mathbf{h}_{2*}}(\theta)=&-\frac{1}{4\pi^{2}}\left(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})}{2}\right)\\ &+\frac{g_{1*}^{2}N}{64\pi^{4}}\left(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*}\frac{\pi-\theta}{\sin\theta}-\frac{(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})}{2}\right)\\ &+\frac{g_{2*}}{384\pi^{4}}\log(64e^{-1})\left(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*}\frac{(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})}{2}\frac{\pi-\theta}{\sin\theta}-\frac{((\mathbf{h}_{1*}^{2})^{2}+(\mathbf{h}_{2*}^{2})^{2})}{2}\right)\\ &-\left(\epsilon\frac{1}{8\pi^{2}}-\frac{g_{1*}^{2}N}{128\pi^{4}}\right)((\mathbf{h}_{1*}\cdot\mathbf{h}_{2*})F^{\prime}_{0}(\cos\theta)-(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2}))\\ &+\frac{g_{2*}}{1536\pi^{8}}\Big((\mathbf{h}_{1*}\cdot\mathbf{h}_{2*})(\mathbf{h}_{1*}^{2}+\mathbf{h}_{2*}^{2})I_{11}(\cos\theta)-(8\pi^{4}-4\pi^{4}\log 8-24\pi^{2}\zeta(3))((\mathbf{h}_{1*}^{2})^{2}+(\mathbf{h}_{2*}^{2})^{2})\Big)\\ &+\frac{g_{2*}}{1024\pi^{8}}\Big(\left(\frac{\mathbf{h}_{1*}^{2}\mathbf{h}_{2*}^{2}+2(\mathbf{h}_{1*}\cdot\mathbf{h}_{2*})^{2}}{3}\right)I_{12}(\cos\theta)-12\pi^{2}\zeta(3)\big((\mathbf{h}_{1*}^{2})^{2}+(\mathbf{h}_{2*}^{2})^{2}\big)\Big)\end{split}$ (2.46)
@@
$\begin{split}\Gamma_{\mathbf{h}_{1}\mathbf{h}_{2}}(\theta)=&-\frac{h_{*}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\\ &+\frac{h_{*}^{2}}{64\pi^{4}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\left(g_{1*}^{2}N+\frac{g_{2*}h_{*}^{2}}{6}\log(64e^{-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{8\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{128\pi^{4}}\right)((\hat{m}_{1}\cdot\hat{m}_{2})F^{\prime}_{0}(\cos\theta)-2)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}((\hat{m}_{1}\cdot\hat{m}_{2})I_{11}(\cos\theta)-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi^{8}}(\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)I_{12}(\cos\theta)-24\pi^{2}\zeta(3))\end{split}$ (2.47)
@@
$\begin{split}\Gamma_{1}(\theta)=&-\frac{h_{*,1}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\\ &-\frac{h_{*,0}^{2}}{4\pi^{2}}\Bigg[-\frac{(N-4+N_{s}+(4-N_{s})\log 64)}{N+8-2N_{s}}\left(\hat{m}_{1}\cdot\hat{m}_{2}\frac{\pi-\theta}{\sin\theta}-1\right)\\ &+\Big(\frac{(8-2N_{s})}{2(N+8-2N_{s})}(\hat{m}_{1}\cdot\hat{m}_{2}F^{\prime}_{0}(\cos\theta)-2)\Big)\\ &-\frac{3(4-N_{s})}{2\pi^{4}(N+8-2N_{s})}\times\\ &\Big(\frac{1}{3}((\hat{m}_{1}\cdot\hat{m}_{2})I_{11}(\cos\theta)-8\pi^{4}+4\pi^{4}\log 8+24\pi^{2}\zeta(3))\\ &+\frac{1}{4}(\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)I_{12}(\cos\theta)-24\pi^{2}\zeta(3))\Big)\Bigg]\end{split}$ (2.50)
@@ 2.3 Large $N$ limit
Figure 2.3: Diagram contributing to the cusp anomalous dimension to leading order at large $N$.
@@ 3 Results and Discussion
Figure 3.1: The term of $\mathcal{O}(\epsilon)$ of the cusp anomalous dimension in the two types of cusped defect in the GNY model.
@@
Figure 3.2: The term of $\mathcal{O}(\epsilon^{0})$ of the cusp anomalous dimension in the two types of cusped defect in the GNY model.
Figure 3.3: Padé Approximation of $\Gamma_{++}(\theta)$ in the $d=3$ GNY model.
@@
$\Delta_{\hat{\phi}}$ (2-loop Padé(1,1)) 1.3673 1.3386 1.3076
@@ 3.1 Dimension of Defect Changing and Creation (Annihilation) Operators
$\begin{split}\Delta_{\hat{m}_{1},\hat{m}_{2}}=&-\frac{h_{*}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}-1\right)\\ &-\frac{h_{*}^{2}}{4\pi^{2}}\left(\hat{m}_{1}\cdot\hat{m}_{2}-1\right)\left(-\frac{g_{1*}^{2}N}{16\pi^{2}}-\frac{g_{2*}h_{*}^{2}}{96\pi^{2}}\log(64e^{-1})\right)\\ &-\left(\epsilon\frac{h_{*}^{2}}{4\pi^{2}}-\frac{g_{1*}^{2}Nh_{*}^{2}}{64\pi^{2}}\right)((\hat{m}_{1}\cdot\hat{m}_{2})-1)\\ &+\frac{g_{2*}h_{*}^{4}}{768\pi^{8}}((\hat{m}_{1}\cdot\hat{m}_{2})-1)(8\pi^{4}-4\pi^{4}\log 8-24\pi^{2}\zeta(3))\\ &+\frac{g_{2*}h_{*}^{4}}{1024\pi^{8}}(\left(\frac{1+2(\hat{m}_{1}\cdot\hat{m}_{2})^{2}}{3}\right)-1)(24\pi^{2}\zeta(3))\end{split}$ (3.1)
@@
$\begin{split}\Delta_{\hat{m}_{1},\hat{m}_{2}}=&(1-\hat{m}_{1}\cdot\hat{m}_{2})\Big[\frac{h_{*}^{2}}{4\pi^{2}}(1+\epsilon)-\frac{g_{1*}^{2}Nh_{*}^{2}}{32\pi^{4}}\\ &-\frac{g_{2*}h_{*}^{4}}{128\pi^{6}}(\pi^{2}-4\zeta(3))+\frac{g_{2*}h_{*}^{4}\zeta(3)}{64\pi^{6}}(1+\hat{m}_{1}\cdot\hat{m}_{2})\Big]\end{split}$ (3.2)
@@
$\begin{split}\Delta_{\hat{m}_{1}\hat{m}_{2}}^{(1)}=&(1-\hat{m}_{1}\cdot\hat{m}_{2})\Bigg[\frac{h_{*,1}^{2}}{4\pi^{2}}+\frac{h_{*,0}^{2}}{4\pi^{4}}\Bigg(\frac{\pi^{2}(N_{2}-N-4)+6(4-N_{s})(3+\hat{m}_{1}\cdot\hat{m}_{2})\zeta(3)}{(8+N-2N_{s})}\Bigg)\Bigg]\end{split}$ (3.5)
@@
We obtain
$\Delta_{0,\mathbf{h}}=\frac{h_{*}^{2}}{8\pi^{2}}(1+\epsilon)-\frac{g_{1*}^{2}Nh_{*,0}^{2}}{64\pi^{4}}-\frac{g_{2*}h_{*,0}^{4}}{256\pi^{6}}\Big(\pi^{2}-\zeta(3)\Big)+\mathcal{O}(\epsilon^{2})$ (3.8)
@@
$\Delta_{0,\mathbf{h}}^{(0)}=\frac{h_{*,0}^{2}}{8\pi^{2}}=\frac{(4-N_{s})(N_{s}+8)}{2\pi^{2}\Big(8-N-2N_{s}+\sqrt{N^{2}+4(-4+N_{s})^{2}+4N(28+5N_{s})}\Big)}$ (3.10)
$\begin{split}\Delta_{0,\mathbf{h}}^{(1)}=\frac{h_{*,1}^{2}}{8\pi^{2}}+\frac{h_{*,0}^{2}}{8\pi^{4}}\Bigg(\frac{\pi^{2}(N_{s}-N-4)+3(4-N_{s})\zeta(3)}{(8+N-2N_{s})}\Bigg)\end{split}$ (3.11)